# DyRoNet: A Low-Rank Adapter Enhanced Dynamic Routing Network for Streaming Perception

### v1

@@ Abstract
Autonomous driving systems demand real-time, accurate perception to navigate complex environments. Addressing this, we introduce the Dynamic Router Network (DyRoNet), a framework that innovates with low-rank dynamic routing for enhanced streaming perception. By integrating specialized pre-trained branch networks, fine-tuned for various environmental conditions, DyRoNet achieves a balance between latency and precision. Its core feature, the speed router module, intelligently directs input data to the best-suited branch network, optimizing performance. The extensive evaluations reveal that DyRoNet adapts effectively to multiple branch selection strategies, setting a new benchmark in performance across a range of scenarios. DyRoNet not only establishes a new benchmark for streaming perception but also provides valuable engineering insights for future work. More project information is available at https://tastevision.github.io/DyRoNet/.
Abstract
The appendix completes the main paper by providing in-depth research details and extended experimental results. The structure of the appendix is organized as follows:
  1. 1.
    Analysis of Environmental Factors Affecting Streaming Perception: Sec. A
    • Impact of Weather Conditions: Sec. A.1
    • Quantitative Analysis of Objects: Sec. A.2
    • Proportion of Small Objects: Sec. A.3
    • Environmental Speed Dynamics: Sec. A.4
  2. 2.
    Expanded Experimental Results: Sec. B
    • Inference Time: Analysis Sec. B.1
  3. 3.
    Detailed Description of DyRoNet: Sec. C
    • Selection of Pre-trained Model: Sec. C.1
    • Hyperparameter Settings: Sec. C.2
@@ 1 Introduction
The quest for low-latency and high-precision perception in autonomous driving systems is paramount. Traditional object detection algorithms Zou et al. (2023), while robust across various scenarios, often grapple with latency issues due to computational demands, creating a disconnect between algorithmic predictions and actual object locations. This latency not only affects the accuracy of detection but also compromises decision-making processes critical to autonomous driving Chen et al. (2023).
In response to these latency challenges, the advent of streaming perception introduces a novel approach to object detection Hauptmann et al. (2023). Unlike conventional methods that focus on immediate detection, streaming perception aims to mitigate processing delays by predicting future environmental states Li et al. (2020). This forward-looking strategy seeks to bridge the gap between processing times and real-world dynamics, significantly improving the real-time responsiveness and reliability of autonomous driving technologies Muhammad et al. (2020). By prioritizing anticipatory processing, streaming perception represents a paradigm shift towards more adaptive and safer autonomous driving systems.
Figure 1: Schematic representation of DyRoNet’s dynamic adaptation in streaming perception. This figure contrasts DyRoNet’s ability to dynamically select the optimal streaming perception method against the limitations of conventional single-method systems in complex scenarios. For easy understanding, it is recommended to view this diagram in color and at an enlarged scale, highlighting DyRoNet’s efficacy in adapting to varied and challenging autonomous driving environments.
@@
Despite the potential of streaming perception, real-world applications reveal inherent contradictions. The unpredictability of driving environments, influenced by variables such as camera movements, weather conditions, lighting changes Zhu et al. (2024), and small objects, significantly affects perception performance. These factors lead to inconsistencies that challenge the system’s robustness and reliability, underscoring the limitations of using a one-size-fits-all model approach (see Sec. 3.1). The diversity of real-world scenarios necessitates more adaptable and versatile models Guo et al. (2019). Key challenges in streaming perception are:
(1) Adapting to Diverse Scenarios: The complexity and variability of autonomous driving environments necessitate models that can adjust to a broad spectrum of conditions (refer to Fig. 1). Designing algorithms that are both specific to certain conditions and universally applicable is challenging. The goal is to modify models to suit various scenarios without losing their essence, a topic we explore further in Sec. 3.1 and Appendix A.
(2) Balancing Accuracy with Efficiency: Integrating models of different scales to cater to diverse scene complexities requires a delicate balance between accuracy and speed. Larger models, while detailed, may slow down the processing time, whereas smaller models can speed up inference at the cost of accuracy. Finding an optimal balance is crucial. Strategies for achieving this balance, including the use of varied architectural designs to strengthen streaming perception, are discussed in Sec. 3.1 and Sec. 4.3.
Addressing the highlighted challenges—diverse scenario distribution and the performance-efficiency balance—is crucial for advancing autonomous driving technologies, as explored in Sec. 3.1. DyRoNet emerges as our solution, a bespoke framework crafted to refine dynamic routing capabilities within autonomous driving systems. This low-rank enhanced dynamic routing framework is specifically developed for streaming perception, featuring a sophisticated speed router module at its core. This module efficiently navigates incoming data to the appropriate pre-trained branch network, each of which is fine-tuned for exceptional performance under varying environmental conditions. The operational details and strategic significance of DyRoNet are further elaborated in Sec. 3.2, showcasing its key role in enhancing the adaptability and performance of autonomous driving systems. To sum up, the contributions are listed as:
@@ 2 Related Work
This section revisits key developments in streaming perception and dynamic neural networks, highlighting differences from our proposed DyRoNet framework. While existing methods have made progress, limitations persist in addressing real-world autonomous driving complexity.
@@ 3.2 Architecture of DyRoNet
Figure 3: The mean curves of frame differences are depicted here. The four curves correspond to frame sizes of the original frame, $\times200$, $\times100$, and $\times50$. Notably, these curves show distinct fluctuations across different vehicle motion scenarios.
@@
Specifically, for frames at times $t$ and t1t-1, represented as $I_{t}$ and $I_{t-1}$ respectively, the frame difference is computed as $\Delta I_{t}=I_{t}-I_{t-1}$. The architecture of the Router Network, $\mathcal{R}$, is simple yet efficient. It consists of a single convolutional layer followed by a linear layer. The network’s output, denoted as $f^{r}\in\mathbb{R}^{K}$, captures the essence of the environmental dynamics. Based on this output, the index $\sigma$ of the optimal branch for processing the current input frame $I_{t}$ is determined through the following equation:
@@
To address these challenges, we have incorporated the Low-Rank Adapter Hu et al. (2021) into our streaming perception models. Within each model $P_{i}$, initially pre-trained on a dataset, the key components are the convolution kernel and bias matrices, symbolized as $W^{d}_{i}$. The rank of the Low-Rank Adaptation (LoRA) module is defined as $r$, a value significantly smaller than the dimensionality of $W^{d}_{i}$, to ensure efficient adaptation. The update to the weight matrix adheres to a low-rank decomposition form, represented as $W_{d}^{i}+\delta W=W_{d}^{i}+BA$. [^1] Here, $B$ is a matrix in $R^{d\times r}$, and $A$ is in $R^{r\times k}$, ensuring that the rank $r$ remains much smaller than $d$. This adaptation strategy allows for the original weights $W_{d}^{i}$ to remain fixed, while the low-rank components $BA$ are trained and adjusted. The adaptation process is executed through the following projection:
@@ 3.3 Training Details of DyRoNet
where $\mathcal{L}^{sp}$ represents the streaming perception loss, and $\mathcal{L}^{E^{2}}$ denotes the effective and efficient (E2) loss, which supervises branch selection.
@@
Effective and Efficient (E2) Loss. During the training phase, streaming perception loss values from all branches are compiled into a vector $v^{sp}\in\mathbb{R}^{K}$, and inference time costs are aggregated into $v^{time}\in\mathbb{R}^{K}$, with $K$ indicating the total number of branches in DyRoNet. To account for hardware variability, a normalized inference time vector $\hat{v}^{time}=\mathrm{softmax}(v^{time})$ is introduced. This vector is derived using the Softmax function to minimize the influence of hardware discrepancies. The representation for effective and efficient (E2) decision-making is defined as:
@@
where $\mathcal{O}$ denotes one-hot encoding, producing a boolean vector of length $K$, with the value of $1$ at the index representing the estimated optimal branch at that moment. The E2 Loss is then formulated as:
@@
Overall, the process of training DyRoNet involves striking a meticulous balance between the SP loss, which ensures the efficacy of each branch, and the E2 loss, which optimizes efficiency. The primary objective of this training is to develop a model that not only delivers high accuracy in perception tasks but also operates within acceptable latency constraints, which is a critical requirement for real-time applications. This balanced approach enables DyRoNet to adapt dynamically to varying computational resources and environmental conditions, thereby maintaining optimal performance in diverse streaming perception scenarios.
@@ 4 Experiments
Methods Latency (ms) sAP $\uparrow$ sAP50 $\uparrow$ sAP75 $\uparrow$ sAPs $\uparrow$ sAPm $\uparrow$ sAPl $\uparrow$
@@ 4.3 Comparision with SOTA Methods
$\text{DAMO-StreamNet}_{\text{S + M}}$ 36.61 33.22
$\text{DAMO-StreamNet}_{\text{S + L}}$ 35.12 39.60
$\text{DAMO-StreamNet}_{\text{M + L}}$ 37.30 37.61
@@
$\text{DAMO-StreamNet}_{\text{S + M}}$ 34.8 35.5
$\text{DAMO-StreamNet}_{\text{S + L}}$ 31.1 37.8
$\text{DAMO-StreamNet}_{\text{M + L}}$ 37.4 37.8
@@ 4.5 Ablation Study
Router Network. To validate the effectiveness of the Router Network based on frame difference, we conducted comparative experiments using frame difference $\Delta I_{t}$, the current frame $I_{t}$, and the concatenation of the current frame with the previous historical frame $I_{t}+I_{t-1}]$ as input modality of the Router Network. The experimental results are presented in Tab. 5. To control variables, in these experiments, we froze the model bank during training and only trained the Router Network. And only three different choices of StreamYOLO is involved in model bank. It can be obverse that using frame difference as input exhibits better performance than other two types of input modalities (35.0 of $\text{StreamYOLO}_{\text{S + L}}$ and 34.6 of $\text{StreamYOLO}_{\text{M + L}}$). This indicates that utilizing frame differences offers significant advantages in comprehending and characterizing environmental speed. Conversely, it also underscores that employing single frames as input or using multiple frames as input renders the lightweight model bank selection model ineffective.
@@ Acknowledgments
Xiang Huang, Jun-Yan He, and Zhi-Qi Cheng collaborated on designing the idea and fully participated in the experiments. Jun-Yan He and Zhi-Qi Cheng provided detailed revisions to the paper and made suggestions regarding the experiments. The manuscript underwent review by Chengyang Li, Wangmeng Xiang, Baigui Sun, and Xiao Wu. Zhi-Qi Cheng managed the entire project as the corresponding author.

### v2

@@ Abstract
The quest for real-time, accurate environmental perception is pivotal in the evolution of autonomous driving technologies. In response to this challenge, we present DyRoNet, a Dynamic Router Network that innovates by incorporating low-rank dynamic routing to enhance streaming perception. DyRoNet distinguishes itself by seamlessly integrating a diverse array of specialized pre-trained branch networks, each meticulously fine-tuned for specific environmental contingencies, thus facilitating an optimal balance between response latency and detection precision. Central to DyRoNet’s architecture is the Speed Router module, which employs an intelligent routing mechanism to dynamically allocate input data to the most suitable branch network, thereby ensuring enhanced performance adaptability in real-time scenarios. Through comprehensive evaluations, DyRoNet demonstrates superior adaptability and improved performance over existing methods, efficiently catering to a wide variety of environmental conditions and setting new benchmarks in streaming perception accuracy and efficiency. Beyond establishing a paradigm in autonomous driving perception, DyRoNet also offers engineering insights and lays a foundational framework for future advancements in streaming perception. For further information and updates on the project, visit https://tastevision.github.io/DyRoNet/.
@@ 1 Introduction
In autonomous driving, achieving low-latency and high-precision perception stands as a critical challenge. Traditional object detection algorithms Zou et al. (2023), though effective across a variety of scenarios, often face latency issues due to high computational demands. This latency leads to a mismatch between algorithmic predictions and the real-world positions of objects, adversely affecting detection accuracy and the decision-making process crucial for safe autonomous navigation Chen et al. (2023). To address these challenges, streaming perception has emerged as a groundbreaking approach, offering a novel perspective on object detection Li et al. (2020). Diverging from traditional methods that concentrate on immediate detection, streaming perception endeavors to predict future environmental conditions, thus aiming to reduce the lag between computation and real-world dynamics. This approach enhances the real-time responsiveness of autonomous driving systems, marking a leap towards adaptive, efficient, and safer vehicular autonomy Hauptmann et al. (2023); Muhammad et al. (2020).
Figure 1: Illustration of DyRoNet’s dynamic selection mechanism in streaming perception. This diagram showcases DyRoNet’s capability to adaptively choose the most suitable perception strategy, contrasting with the static approach of traditional methods in complex environments [Viewing in color and at an expanded scale].
@@
Despite its advantages, streaming perception’s application in unpredictable and dynamic autonomous driving reveals key challenges. The performance of perception systems is laboriously influenced by factors such as camera motion, weather variability, lighting changes Zhu et al. (2024), and the presence of small objects. These factors introduce performance inconsistencies, challenging the robustness and reliability of a singular, one-size-fits-all model approach (see Sec. 3.1). Such diversity in driving environments mandates the development of more adaptable and versatile perception models Guo et al. (2019):
(1) Adaptation to Diverse Scenarios: The wide range of conditions encountered in autonomous driving requires models capable of adjusting to an extensive variety of scenarios (refer to Fig. 1). Crafting algorithms that maintain their core effectiveness while being adaptable to specific environmental conditions presents a significant challenge, further discussed in Sec. 3.1 and Appendix A.
(2) Balancing Accuracy and Efficiency: Achieving a balance between precision and processing speed when employing models of varying complexities is vital. High-detail models may enhance accuracy but at the cost of increased latency, while simpler models might offer faster inference times but reduced precision. Identifying the optimal compromise between these extremes is essential for enhancing streaming perception’s effectiveness, as Sec. 3.1 and Sec. 4.3.
Addressing the critical challenges of diverse scenario distribution and performance-efficiency balance is paramount for the progression of autonomous driving technologies, a theme extensively discussed in Sec. 3.1. In response, we introduce DyRoNet, a tailored framework designed to enhance the dynamic routing capabilities essential for autonomous systems. This framework, developed specifically for streaming perception, integrates a low-rank dynamic routing mechanism enriched with a sophisticated speed router module. This central module adeptly directs incoming data to the most suitable pre-trained branch network, with each branch meticulously optimized for peak performance across a spectrum of environmental conditions. Detailed operational insights and the strategic underpinnings of DyRoNet are elaborated upon in Sec. 3.2, highlighting its key contribution to augmenting the adaptability and efficiency of autonomous driving systems. The main contributions can be summarized as:
@@ 2 Related Work
This section revisits developments in streaming perception and dynamic neural networks, highlighting differences from our proposed DyRoNet framework. While existing methods have made progress, limitations persist in addressing real-world autonomous driving complexity.
@@ 3.2 Architecture of DyRoNet
Figure 3: The mean curves of frame differences are depicted here. The four curves correspond to frame sizes of the original frame, 200$\times$200, 100$\times$100, and 50$\times$50. Notably, these curves show distinct fluctuations across different vehicle motion scenarios.
@@
Specifically, for frames at times $t$ and $t-1$, represented as $I_{t}$ and $I_{t-1}$ respectively, the frame difference is computed as $\Delta I_{t}=I_{t}-I_{t-1}$. The architecture of the Router Network, $\mathcal{R}$, is simple yet efficient. It consists of a single convolutional layer followed by a linear layer. The network’s output, denoted as $f^{r}\in\mathbb{R}^{K}$, captures the essence of the environmental dynamics. Based on this output, the index $\sigma$ of the optimal branch for processing the current input frame $I_{t}$ is determined through the following equation:
@@
To address these challenges, we have incorporated the Low-Rank Adapter Hu et al. (2021) into our streaming perception models. Within each model $P_{i}$, initially pre-trained on a dataset, the key components are the convolution kernel and bias matrices, symbolized as $W^{d}_{i}$. The rank of the Low-Rank Adaptation (LoRA) module is defined as $r$, a value significantly smaller than the dimensionality of $W^{d}_{i}$, to ensure efficient adaptation. The update to the weight matrix adheres to a low-rank decomposition form, represented as $W_{d}^{i}+\delta W=W_{d}^{i}+BA$ [^1] Here, $B$ is a matrix in $R^{d\times r}$, and $A$ is in $R^{r\times k}$, ensuring that the rank $r$ remains much smaller than $d$. This adaptation strategy allows for the original weights $W_{d}^{i}$ to remain fixed, while the low-rank components $BA$ are trained and adjusted. The adaptation process is executed through the following projection:
@@ 3.3 Training Details of DyRoNet
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 4 Experiments
[…]
@@ 4.3 Comparision with SOTA Methods
[…]
@@
[…]
@@ 4.5 Ablation Study
Router Network. To validate the effectiveness of the Router Network based on frame difference, we conducted comparative experiments using frame difference $\Delta I_{t}$, the current frame $I_{t}$, and the concatenation of the current frame with the previous historical frame $[I_{t}+I_{t-1}]$ as input modality of the Router Network. The experimental results are presented in Tab. 5. To control variables, in these experiments, we froze the model bank during training and only trained the Router Network. And only three different choices of StreamYOLO is involved in model bank. It can be obverse that using frame difference as input exhibits better performance than other two types of input modalities (35.0 of $\text{StreamYOLO}_{\text{S + L}}$ and 34.6 of $\text{StreamYOLO}_{\text{M + L}}$). This indicates that utilizing frame differences offers significant advantages in comprehending and characterizing environmental speed. Conversely, it also underscores that employing single frames as input or using multiple frames as input renders the lightweight model bank selection model ineffective.
@@ Acknowledgments
Xiang Huang, Jun-Yan He, and Zhi-Qi Cheng worked together on the research idea and were fully involved in the experimental work. Jun-Yan He and Zhi-Qi Cheng made revisions to the manuscript, offering valuable insights and recommendations for the experimental procedures. The manuscript was then reviewed by Chengyang Li, Wangmeng Xiang, Baigui Sun, and Xiao Wu. Zhi-Qi Cheng served as the corresponding author and led the entire project.